Please help me answer this word problem:
The rate at which a radioactive isotope disintegrates is proportional to the amount present. If a 30g sample will contain only 20g after ten minutes, what is the half-life of this isotope, correct to the nearest tenth of a minute? Answer this question by first finding the amount y of this isotope as a function of time t(in minutes) in the form y=y0e^kt.
Thank-you, your help is always appreciated!

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

The rate at which a radioactive isotope disintegrates is proportional to the amount present. If a 30g sample will contain only 20g after ten minutes, what is the half-life of this isotope, correct to the nearest tenth of a minute? Answer this question by first finding the amount y of this isotope as a function of time t(in minutes) in the form
y = yo*e^(kt)
------
20 = 30*e^(k*(1/6))
----
e^(k/6) = 2/3
-----
k/6 = ln(2/3) = -0.4055
-----
k = -2.4328
------
Substitute for "k" in the formula:
y = yo*e^(-2.4328t)
---
Note: Half-life means y = yo/2
-----
Then e^(-2.4328t) = 1/2
----
-2.4328t = ln(1/2) = -0.6932
----
time = -0.6932/-2.4328 = 0.2849 hrs = 17min 5.7sec
----------
Cheers,
Stan H.

Answer by solver91311(24713)





Substitute the values you know:



So



Take the log of both sides





Now that you know the value of , solve for in





Take the log



Substitute value of and solve for



The rest is just calculator work.

John

My calculator said it, I believe it, that settles it